OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. x + 2y – z = 2 3x + y – 2z = –15 3x + 4y – 3z = –2

OpenStudy (anonymous):

Eq. 1: x + 2y - z = 2 z = x + 2y - 2 Substitute into Eq. 2 3x + y - 2(x + 2y - 2) = -15 x - 3y + 4 = -15 x = 3y - 19 Substitute values for x and z into Eq. 3 3(3y - 19) + 4y - 3((3y - 19) + 2y - 2) = -2 9y - 57 + 4y - 15y + 63 = -2 -2y + 6 = -2 2y = 8 y = 4 And then you repeat the whole process again. A bit tedious but it's just manipulating algebraic equations...

OpenStudy (anonymous):

thank you.